Use a half angle identity to find the exact value of cos(19pi/12)
Answer by stanbon(75887)

Use a half angle identity to find the exact value of cos(19pi/12)
-------------------
Note: cos(t/2) = +/-sqrt[(1+cos(t))/2]
If t/2 = 19pi/12, t = 19pi/6
----
But 19pi/6 = 2pi + (7/6)pi
And cos(7pi/6) = -cos(pi/6) = -sqrt(3)/2
----
So cos(19pi/12) = sqrt[(1-sqrt(3)/2))/2] = sqrt(2-sqrt(3))/4)
= sqrt(2-sqrt(3))/2
------------------
Cheers,
Stan H.